OpenStudy (anonymous):

find the function f such that f'(x)=1/x^2+1 and f(0)=2

OpenStudy (anonymous):

F = -1/x+x+C I integrated the function And then I plugged in f(0)=2: -1/(0)+(0)+C=2 C=2 F=x-(1/x)+2

OpenStudy (anonymous):

i thought the derivative of 1/x^2+1 is arctan

OpenStudy (anonymous):

yup that is an identity ugghhhh

OpenStudy (anonymous):

wait let me redo it

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

so when that function is integrated you get arctan(x)+c So we tehn plug in f(0)=2: tan^-1(0)+C=2 0+C=2 C=2 F=tan^-1(x)+2

OpenStudy (anonymous):

thank u